The Frontive Company has assembled the following data pertaining to certain costs that cannot be easily...

  1. The Frontive Company has assembled the following data pertaining to certain costs that cannot be easily identified as either fixed or variable. Frontive Company has heard about a method of measuring cost functions called the high-low method and has decided to use it in this situation.

Cost

Hours

$24,700

5,000

26,000

5,500

34,000

7,500

45,370

10,300

38,000

9,500

What is the estimated total cost at an operating level of 9,000 hours?

  1. $40,300
  2. $39,644
  3. $44,460
  4. $35,100
  5. None of the above
Answer #1

Variable cost per unit=[Total cost at highest level-Total cost at lowest level]/(Highest level-Lowest level)

=(45370-24700)/(10300-5000)

=$3.9 per hour

Hence Total fixed cost=45370-(3.9*10300)=5200

Hence estimated total cost=Total variable cost+Total fixed cost

=(3.9*9000)+5200

=$40300